# Self-Improving Skill Learning for Robust Skill-based Meta-Reinforcement Learning

### v3

@@ Abstract
Meta-reinforcement learning (Meta-RL) facilitates rapid adaptation to unseen tasks but faces challenges in long-horizon environments. Skill-based approaches tackle this by decomposing state-action sequences into reusable skills and employing hierarchical decision-making. However, these methods are highly susceptible to noisy offline demonstrations, leading to unstable skill learning and degraded performance. To address this, we propose Self-Improving Skill Learning (SISL), which performs self-guided skill refinement using decoupled high-level and skill improvement policies, while applying skill prioritization via maximum return relabeling to focus updates on task-relevant trajectories, resulting in robust and stable adaptation even under noisy and suboptimal data. By mitigating the effect of noise, SISL achieves reliable skill learning and consistently outperforms other skill-based meta-RL methods on diverse long-horizon tasks.
@@ 1 Introduction
Reinforcement Learning (RL) has achieved significant success in domains such as game environments and robotic control (Mnih et al., 2015; Andrychowicz et al., 2020). However, it struggles to adapt quickly to new tasks. Meta-RL addresses this limitation by enabling rapid adaptation to unseen tasks through meta-learning how policies solve problems (Duan et al., 2016; Finn et al., 2017). Among various approaches, context-based meta-RL stands out for its ability to represent similar tasks with analogous contexts and leverage this information in the policy, facilitating quick adaptation to new tasks (Rakelly et al., 2019; Zintgraf et al., 2019). Notably, PEARL (Rakelly et al., 2019) has been widely studied for its high sample efficiency, achieved through off-policy learning, which allows for the reuse of previous samples. Despite these strengths, existing meta-RL methods face challenges in long-horizon environments, where extracting meaningful context information becomes difficult, hindering effective learning.
@@
Skill-based approaches address these challenges by breaking down long state-action sequences into reusable skills, facilitating hierarchical decision-making and enhancing efficiency in complex tasks (Pertsch et al., 2021; 2022; Shi et al., 2023). Among these, SPiRL (Pertsch et al., 2021) defines skills as temporal abstractions of actions, employing them as low-level policies within a hierarchical framework to achieve success in long-horizon tasks. SiMPL (Nam et al., 2022) builds on this by extending skill learning to meta-RL, using offline expert data to train skills and a context-based high-level policy for task-specific skill selection. Despite these advancements, such methods are highly susceptible to noisy offline demonstrations, which can destabilize skill learning and reduce reliability. In real-world settings, noise often stems from factors like infrastructure aging or environmental perturbations, making it crucial to design methods that remain robust under such conditions (Brys et al., 2015; Chae et al., 2022; Yu et al., 2024a).
@@
While noisy demonstration handling has been explored in other RL settings (Sasaki & Yamashina, 2020; Mandlekar et al., 2022), skill-based meta-RL has largely overlooked this challenge. We identify a critical failure mode: when offline data are suboptimal, the skill library becomes corrupted, and this degradation propagates to the high-level policy, ultimately harming adaptation performance. To address this, we propose Self-Improving Skill Learning (SISL), a robust skill-based meta-RL framework with two key contributions: (1) Decoupled skill self-improvement, achieved through a dedicated improvement policy that perturbs trajectories near the offline data distribution, discovers higher-quality rollouts, and supervises its own updates via a prioritized online buffer. This process progressively denoises the skill library while preserving stability. (2) Skill prioritization via maximum return relabeling, which evaluates offline trajectories with a learned reward model, assigns task-relevant hypothetical returns, and reweights them through a softmax prioritization scheme. This suppresses noisy or irrelevant samples and focuses skill updates on the most beneficial trajectories for downstream adaptation. Together, these components dynamically balance offline and online data contributions, yielding a progressively cleaner skill library and accelerating meta-RL convergence. To our knowledge, SISL is the first framework to explicitly address suboptimal demonstrations in skill-based meta-RL through both exploration-guided refinement and principled relabeling, significantly improving robustness and generalization of skill-learning in real-world noisy scenarios.
@@ 2 Related Works
Skill-based Reinforcement Learning: Skill-based RL has gained traction for tackling complex tasks by leveraging temporally extended actions. Researchers have proposed information-theoretic approaches to discover diverse and predictable skills (Gregor et al., 2016; Eysenbach et al., 2018; Achiam et al., 2018; Sharma et al., 2019), with recent work improving skill quality through additional constraints and objectives (Strouse et al., 2022; Park et al., 2022; 2023; Hu et al., 2024). In offline scenarios, approaches focus on learning transferable behavior priors and hierarchical skills from demonstration data (Pertsch et al., 2021; 2022; Shi et al., 2023; Xu et al., 2022; Kipf et al., 2019; Rana et al., 2023). Building upon these foundations, various skill-based meta-RL approaches have been developed, from hierarchical and embedding-based methods (Nam et al., 2022; Chien & Lai, 2023; Cho & Sun, 2024) to task decomposition strategies (Yoo et al., 2022; He et al., 2024) and unsupervised frameworks (Gupta et al., 2018; Jabri et al., 2019; Shin et al., 2024). For clarity, we use skill in the canonical skill-based RL sense: a latent over fixed-length action sequences. Therefore, studies (Yu et al., 2024b; Wu et al., ) that lack explicit skill learning are not categorized here.
@@
Relabeling Techniques for Meta-RL: Recent developments in meta-RL have introduced various relabeling techniques to enhance sample efficiency and task generalization (Pong et al., 2022; Jiang et al., 2023). Goal relabeling approaches have extended hindsight experience replay to meta-learning contexts (Packer et al., 2021; Wan et al., 2021), enabling agents to learn from failed attempts. For reward relabeling, model-based approaches have been proposed to relabel experiences across different tasks (Mendonca et al., 2020), improving adaptation to out-of-distribution scenarios. Beyond these categories, some methods have introduced innovative relabeling strategies using contrastive learning (Yuan & Lu, 2022; Zhou et al., 2024) and metric-based approaches (Li et al., 2020) to create robust task representations in offline settings.
@@
Hierarchical Frameworks: Hierarchical approaches in RL have been pivotal for solving long-horizon tasks, where various methods have been proposed including goal-conditioned learning (Levy et al., 2019; Li et al., 2019; Gehring et al., 2021), and option-based frameworks (Bacon et al., 2017; Riemer et al., 2018; Barreto et al., 2019; Araki et al., 2021). The integration of hierarchical frameworks with meta-RL has shown significant potential for rapid task adaptation and complexity handling (Frans et al., 2018; Fu et al., 2020a; 2023). Recent work has demonstrated that hierarchical architectures in meta-RL can provide theoretical guarantees for learning optimal policies (Chua et al., 2023) and achieve efficient learning through transformer-based architectures (Shala et al., 2024). Recent advances in goal-conditioned RL have focused on improving sample efficiency (Robert et al., 2024), state representation (Yin et al., 2024), and offline-to-online RL (Park et al., 2024; Schmidt et al., 2025). Offline-to-online RL assumes reward-annotated offline datasets to pretrain the policy based on RL before online fine-tuning. In contrast, our setting provides only reward-free offline data for skill learning, making direct application of offline-to-online RL infeasible and clearly distinguishing our approach.
@@ 3 Background
where $\mathcal{L}(\pi_{l},q,p,z):=-\sum^{t+H_{s}-1}_{k=t}\log\pi_{l}(a_{k}|s_{k},z)+\lambda^{\text{kld}}_{l}\mathcal{D}_{\text{KL}}\left(q||\mathcal{N}(\mathbf{0},\mathbf{I})\right)+\mathcal{D}_{\text{KL}}(\lfloor q\rfloor||p)$, $H_{s}$ is the skill length, $\lambda^{\text{kld}}_{l}$ is the coefficient for KL divergence (KLD) $\mathcal{D}_{\text{KL}}$, $\lfloor\cdot\rfloor$ is the stop gradient operator, and $\mathcal{N}(\bm{\mu},\mathbf{\Sigma})$ represents a Normal distribution with mean $\bm{\mu}$ and covariance matrix $\mathbf{I}$. Here, $p(z|s_{t})$ is the skill prior to obtain the skill distribution $z$ for a given state $s_{t}$ directly. Using the learned skill policy $\pi_{l}$, the high-level policy $\pi_{h}$ is trained within a hierarchical framework using RL methods.
@@ 4.1 Motivation: Toward Robust Skill Learning under Noisy Demonstrations
Figure 2: Comparison of prior skill learning methods in microwave-opening task: (a) Learned skills with expert and noisy demonstrations. (b) Meta-RL performance with the learned skills.
@@ 4.2 Self-Improving Skill Refinement with Decoupled Policies
$\sum_{i}\mathbb{E}_{\tau^{i}\sim\mathcal{B}_{\mathrm{imp}}^{i}\cup\mathcal{B}_{\mathrm{on}}^{i}}\big[\mathcal{L}^{\mathrm{RL}}_{\mathrm{imp}}(\pi_{\mathrm{imp}})\big]+\lambda^{\mathrm{kld}}_{\mathrm{imp}}\mathbb{E}_{\tau^{i}\sim\mathcal{B}_{\mathrm{on}}^{i}}\mathcal{D}_{\mathrm{KL}}(\hat{\pi}_{d}^{i}\|\pi_{\mathrm{imp}}),\vskip-5.0pt$ (3)
@@
Building on the improved trajectory distribution obtained from this update, SISL performs skill refinement in dedicated phases every $K_{\mathrm{iter}}$ iterations rather than after every update step. At the end of each phase, the skill encoder $q$, skill prior $p$, and low-level policy $\pi_{l}$ are re-trained on both $\mathcal{B}_{\mathrm{off}}$ and $\mathcal{B}_{\mathrm{on}}^{i}$ using the SPiRL objective in Eq. 1, and the high-level policy $\pi_{h}$ is reinitialized to fully benefit from the updated skill library. This periodic refinement mitigates bias from outdated skill embeddings and accelerates adaptation, as shown in our ablation study in Appendix G.4. These phases yield progressively denoised skills and a stable training signal, enabling $\pi_{h}$ to solve tasks more efficiently as training progresses. The overall SISL structure is illustrated in Fig. 4. However, the large number of noisy trajectories in $\mathcal{B}_{\mathrm{off}}$ can still reduce skill learning efficiency. The next section introduces a maximum return relabeling mechanism that prioritizes the most relevant offline trajectories so that only high-value samples significantly influence skill refinement.
@@ 4.3 Skill Prioritization via Maximum Return Relabeling
Figure 3: The SISL framework
Figure 4: Illustration of maximum return relabeling
@@
where $\mathcal{L}(\pi_{l},q,p,z)$ is the SPiRL skill loss defined in Eq. 1. Since $\mathcal{B}_{\text{on}}$ already contains high-return trajectories, samples are drawn uniformly from this buffer. In addtion, the mixing coefficient $\beta$ is computed dynamically from the offline and online datasets based on their average returns, adaptively balancing their contributions during training as
@@ 5 Experiment
In this section, we evaluate the robustness of the SISL framework to noisy demonstrations in long-horizon environments and analyze how self-improving skill learning enhance performance.
@@ 5.1 Experimental Setup
We compare the proposed SISL with 3 non-meta RL baselines: SAC, which trains test tasks directly without using the offline dataset; SAC+RND, which incorporates RND-based intrinsic noise for enhanced exploration; and SPiRL, which learns skills from the offline dataset using Eq. (1) and trains high-level policies for individual tasks. Also, we include 4 meta-RL baselines: PEARL, a widely used context-based meta-RL algorithm without skill learning; PEARL+RND, which integrates RND-based exploration into PEARL; SiMPL, which applies skill-based meta-RL using Eq. (2); and our SISL. SISL’s hyperparameters primarily follow Nam et al. (2022), with additional parameters (e.g., temperature $T$) tuned via hyperparameter search, while other baselines use author-provided code. Results are averaged over 5 random seeds, with standard deviations represented as shaded areas in graphs and $\pm$ values in tables.
5.2 Environmental Setup

### v4

@@ Abstract
Meta-reinforcement learning (Meta-RL) facilitates rapid adaptation to unseen tasks but faces challenges in long-horizon environments. Skill-based approaches tackle this by decomposing state-action sequences into reusable skills and employing hierarchical decision-making. However, these methods are highly susceptible to noisy offline demonstrations, leading to unstable skill learning and degraded performance. To address this, we propose Self-Improving Skill Learning (SISL), which performs self-guided skill refinement using decoupled high-level and skill improvement policies, while applying skill prioritization via maximum return relabeling to focus updates on task-relevant trajectories, resulting in robust and stable adaptation even under noisy and suboptimal data. By mitigating the effect of noise, SISL achieves reliable skill learning and consistently outperforms other skill-based meta-RL methods on diverse long-horizon tasks. Our code is available at https://github.com/epsilog/SISL.
@@ 1 Introduction
Reinforcement Learning (RL) has achieved significant success in various domains (Mnih et al., 2015; Andrychowicz et al., 2020). While advanced exploration frameworks have been proposed to improve sample efficiency (Pathak et al., 2017; Fortunato et al., 2018; Han and Sung, 2021b; a; Jo et al., 2024), traditional RL struggles to adapt quickly to new tasks. Meta-RL addresses this limitation by enabling rapid adaptation to unseen tasks through meta-learning how policies solve problems (Duan et al., 2016; Finn et al., 2017). Among various approaches, context-based meta-RL stands out for its ability to represent similar tasks with analogous contexts and leverage this information in the policy, facilitating quick adaptation to new tasks (Rakelly et al., 2019; Zintgraf et al., 2019; Kim et al., 2025). Notably, PEARL (Rakelly et al., 2019) has been widely studied for its high sample efficiency, achieved through off-policy learning, which allows for the reuse of previous samples. Despite these strengths, existing meta-RL methods face challenges in long-horizon environments, where extracting meaningful context information becomes difficult, hindering effective learning.
@@
Skill-based approaches address these challenges by breaking down long state-action sequences into reusable skills, facilitating hierarchical decision-making and enhancing efficiency in complex tasks (Pertsch et al., 2021; 2022; Shi et al., 2023). Among these, SPiRL (Pertsch et al., 2021) defines skills as temporal abstractions of actions, employing them as low-level policies within a hierarchical framework to achieve success in long-horizon tasks. SiMPL (Nam et al., 2022) builds on this by extending skill learning to meta-RL, using offline expert data to train skills and a context-based high-level policy for task-specific skill selection. Despite these advancements, such methods are highly susceptible to noisy offline demonstrations, which can destabilize skill learning and reduce reliability. In real-world settings, noise often stems from factors like infrastructure aging or environmental perturbations, making it crucial to design methods that remain robust under such conditions (Brys et al., 2015; Chae et al., 2022; Yu et al., 2024a; Lee et al., 2025).
@@
While noisy demonstration handling has been explored in other RL settings (Sasaki and Yamashina, 2020; Mandlekar et al., 2022), skill-based meta-RL has largely overlooked this challenge. We identify a critical failure mode: when offline data are suboptimal, the skill library becomes corrupted, and this degradation propagates to the high-level policy, ultimately harming adaptation performance. To address this, we propose Self-Improving Skill Learning (SISL), a robust skill-based meta-RL framework with two key contributions: (1) Decoupled skill self-improvement, using an improvement policy that perturbs trajectories near the offline data distribution, discovers higher-quality rollouts, and self-supervises updates via a prioritized online buffer. This process progressively denoises the skill library while preserving stability. (2) Skill prioritization via maximum return relabeling, which evaluates offline trajectories with a learned reward model, assigns task-relevant hypothetical returns, and reweights them through a softmax prioritization scheme. This suppresses noisy or irrelevant samples and focuses skill updates on the most beneficial trajectories for adaptation. These components dynamically balance offline and online data contributions, yielding a progressively cleaner skill library and accelerating meta-RL convergence. To our knowledge, SISL is the first framework to explicitly address suboptimal demonstrations in skill-based meta-RL, significantly improving robustness and generalization of skill-learning in real-world noisy scenarios.
@@ 2 Related Works
Skill-based Reinforcement Learning: Skill-based RL has gained traction for tackling complex tasks by leveraging temporally extended actions. Researchers have proposed information-theoretic approaches to discover diverse and predictable skills (Gregor et al., 2016; Eysenbach et al., 2018; Achiam et al., 2018; Sharma et al., 2019), with recent work improving skill quality through additional constraints and objectives (Strouse et al., 2022; Park et al., 2022; 2023; Hu et al., 2024). In offline scenarios, approaches focus on learning transferable behavior priors and hierarchical skills from demonstration data (Pertsch et al., 2021; 2022; Shi et al., 2023; Xu et al., 2022; Kipf et al., 2019; Rana et al., 2023). Building upon these foundations, various skill-based meta-RL approaches have been developed, from hierarchical and embedding-based methods (Nam et al., 2022; Chien and Lai, 2023; Cho and Sun, 2024) to task decomposition strategies (Yoo et al., 2022; He et al., 2024) and unsupervised frameworks (Gupta et al., 2018; Jabri et al., 2019; Shin et al., 2024). For clarity, we use skill in the canonical skill-based RL sense: a latent over fixed-length action sequences. Therefore, studies (Yu et al., 2024b; Wu et al., 2019) that lack explicit skill learning are not categorized here.
@@
Relabeling Techniques for Meta-RL: Recent developments in meta-RL have introduced various relabeling techniques to enhance sample efficiency and task generalization (Pong et al., 2022; Jiang et al., 2023). Goal relabeling approaches have extended hindsight experience replay to meta-learning contexts (Packer et al., 2021; Wan et al., 2021), enabling agents to learn from failed attempts. For reward relabeling, model-based approaches have been proposed to relabel experiences across different tasks (Mendonca et al., 2020), improving adaptation to out-of-distribution scenarios. Beyond these categories, some methods have introduced innovative relabeling strategies using contrastive learning (Yuan and Lu, 2022; Zhou et al., 2024) and metric-based approaches (Li et al., 2020) to create robust task representations in offline settings.
@@
Hierarchical Frameworks: Hierarchical approaches in RL have been pivotal for solving long-horizon tasks, including goal-conditioned learning (Levy et al., 2019; Li et al., 2019; Gehring et al., 2021), and option-based frameworks (Bacon et al., 2017; Riemer et al., 2018; Barreto et al., 2019; Araki et al., 2021). The integration of hierarchical frameworks with meta-RL has shown potential for rapid task adaptation and complexity handling (Frans et al., 2018; Fu et al., 2020a; 2023). Recent work has demonstrated that hierarchical architectures in meta-RL can provide theoretical guarantees for learning optimal policies (Chua et al., 2023) and achieve efficient learning through transformer-based architectures (Shala et al., 2024). Recent advances in goal-conditioned RL have focused on improving sample efficiency (Robert et al., 2024; Hwang et al., 2025), state representation (Yin et al., 2024), and offline-to-online RL (Park et al., 2024; Schmidt et al., 2025). Offline-to-online RL assumes reward-annotated offline datasets to pretrain the policy based on RL before online fine-tuning. In contrast, our setting provides only reward-free offline data for skill learning, making direct application of offline-to-online RL infeasible and clearly distinguishing our approach.
@@ 3 Background
where $\mathcal{L}(\pi_{l},q,p,z):=-\sum^{t+H_{s}-1}_{k=t}\log\pi_{l}(a_{k}|s_{k},z)+\lambda^{\text{kld}}_{l}\mathcal{D}_{\text{KL}}\left(q||\mathcal{N}(\mathbf{0},\mathbf{I})\right)+\mathcal{D}_{\text{KL}}(\lfloor q\rfloor||p)$, $H_{s}$ is the skill length, $\lambda^{\text{kld}}_{l}$ is the coefficient for KL divergence (KLD) $\mathcal{D}_{\text{KL}}$, $\lfloor\cdot\rfloor$ is the stop gradient operator, and $\mathcal{N}(\bm{\mu},\mathbf{\Sigma})$ represents a Normal distribution with mean $\bm{\mu}$ and covariance matrix $\mathbf{\Sigma}$. Here, $p(z|s_{t})$ is the skill prior to obtain the skill distribution $z$ for a given state $s_{t}$ directly. Using the learned skill policy $\pi_{l}$, the high-level policy $\pi_{h}$ is trained within a hierarchical framework using RL methods. In our paper, the skill refinement procedure builds on Eq. (1) and is applied during both initial skill learning phase and subsequent refinement. A more detailed description is provided in Appendix B.1.
@@ 4.1 Motivation: Toward Robust Skill Learning under Noisy Demonstrations
Figure 2: Comparison of prior skill learning method in microwave-opening task: (a) Learned skills with expert and noisy demonstrations. (b) Meta-RL performance with learned skills. Low-level skills are learned via SPiRL, while the meta-learning component follows the structure of SiMPL.
@@ 4.2 Self-Improving Skill Refinement with Decoupled Policies
Figure 3: Proposed SISL framework. The decoupled policy learning phase uses ($\pi_{h}$, $\pi_{\text{imp}}$) to solve tasks and discover improved behaviors. The skill learning phase periodically updates the skill components $(\pi_{l},p,q)$.
Figure 4: Maximum return relabeling. After relabeling to the online buffers $\mathcal{B}_{\text{on}}^{i}$ and the offline buffer $\mathcal{B}_{\text{off}}$ using the learned reward model, SISL performs prioritized skill refinement by mixing both buffers via coefficient $\beta$.
@@
$\sum_{i}\mathbb{E}_{\tau^{i}\sim\mathcal{B}_{\mathrm{imp}}^{i}\cup\mathcal{B}_{\mathrm{on}}^{i}}\big[\mathcal{L}^{\mathrm{RL}}_{\mathrm{imp}}(\pi_{\mathrm{imp}})\big]+\lambda^{\mathrm{kld}}_{\mathrm{imp}}\mathbb{E}_{\tau^{i}\sim\mathcal{B}_{\mathrm{on}}^{i}}\mathcal{D}_{\mathrm{KL}}(\hat{\pi}_{d}^{i}\|\pi_{\mathrm{imp}}),$ (3)
@@
Building on the improved trajectory distribution obtained from this update, SISL performs skill refinement in dedicated phases every $K_{\mathrm{iter}}$ iterations rather than after every update step. At the end of each phase, the skill encoder $q$, skill prior $p$, and low-level policy $\pi_{l}$ are re-trained on both $\mathcal{B}_{\mathrm{off}}$ and $\mathcal{B}_{\mathrm{on}}^{i}$ using the SPiRL objective in Eq. (1), and the high-level policy $\pi_{h}$ is reinitialized to fully benefit from the updated skill library. This periodic refinement mitigates bias from outdated skill embeddings and accelerates adaptation, as shown in our ablation study in Appendix G.4. These phases yield progressively denoised skills and a stable training signal, enabling $\pi_{h}$ to solve tasks more efficiently as training progresses. The overall SISL structure is illustrated in Fig. 4. However, the large number of noisy trajectories in $\mathcal{B}_{\mathrm{off}}$ can still reduce skill learning efficiency. The next section introduces a maximum return relabeling mechanism that prioritizes the most relevant offline trajectories so that only high-value samples significantly influence skill refinement.
@@ 4.3 Skill Prioritization via Maximum Return Relabeling
where $\mathcal{L}(\pi_{l},q,p,z)$ is the SPiRL skill loss defined in Eq. (1). Since $\mathcal{B}_{\text{on}}$ already contains high-return trajectories, samples are drawn uniformly from this buffer. In addition, the mixing coefficient $\beta$ is computed dynamically from the offline and online datasets based on their average returns, adaptively balancing their contributions during training as
@@ 5 Experiment
In this section, we evaluate the robustness of the SISL framework to noisy demonstrations in long-horizon environments and analyze how self-improving skill learning enhances performance.
@@ 5.1 Experimental Setup
We compare the proposed SISL with 3 non-meta RL baselines: SAC, which trains test tasks directly without using the offline dataset; SAC+RND, which incorporates RND-based intrinsic noise for enhanced exploration; and SPiRL, which learns skills from the offline dataset using Eq. (1) and trains high-level policies for individual tasks. Also, we include 4 meta-RL baselines: PEARL, a widely used context-based meta-RL algorithm without skill learning; PEARL+RND, which integrates RND-based exploration into PEARL; SiMPL, which applies skill-based meta-RL using Eq. (2); and our SISL. SISL’s hyperparameters primarily follow Nam et al. (2022), with additional parameters (e.g., temperature $T$) tuned via hyperparameter search, while other baselines use author-provided code. Although SISL introduces an additional improvement policy $\pi_{\text{imp}}$, we ensured a fair comparison by keeping the total amount of training and interaction the same. Specifically, in each iteration only half of the sampled meta-train tasks use $\pi_{h}$ and the other half use $\pi_{\text{imp}}$, so the total number of samples and policy updates does not exceed that of SiMPL. Results are averaged over 5 random seeds, with standard deviations represented as shaded areas in graphs and $\pm$ values in tables.
@@
5.2 Environmental Setup
5.3 Performance Comparison
We compare SISL with various baseline algorithms. Non-meta RL algorithms are trained directly on each test task for 0.5K iterations due to the absence of a meta-train phase. Meta-RL algorithms undergo meta-training for 10K iterations in Kitchen and Office, and 4K in Maze2D and AntMaze, followed by fine-tuning on test tasks for an additional 0.5K iterations. For SISL, the skill refinement interval $K_{\text{iter}}$ is set to 2K for Kitchen, Office, and AntMaze; 1K for Maze2D. To ensure a fair comparison, SISL counts each update process from its skill-improvement policy and high-level policy as one iteration. Table 1 shows final average return across test tasks after the specified test iterations. In these environments, Kitchen and Office assign a reward of 1 for each successfully completed subtask, while Maze2D and AntMaze give a reward of 1 upon reaching the goal, meaning the return directly corresponds to the success rate. The corresponding learning curves are in Appendix D.1 for a more detailed comparison. From the result, SAC and PEARL baselines, which do not utilize skills or offline datasets, perform poorly on long-horizon tasks, yielding a single result across all noise levels. In contrast, SPiRL, SiMPL, and SISL, which leverage skills, achieve better performance.
SPiRL and SiMPL, however, show sharp performance declines as dataset noise increases. While both perform well with expert data, SiMPL struggles under noisy conditions due to instability in its task encoder $q_{e}$, sometimes performing worse than SPiRL. Here, the baseline results for Maze2D (Expert) are somewhat lower than those reported in the SiMPL paper. This discrepancy likely arises because, in constructing the offline dataset, we considered fewer tasks compared to SiMPL, resulting in trajectories that do not fully cover the map. Interestingly, minor noise occasionally boosts performance by introducing diverse trajectories that improve skill learning, a detailed analysis of changes in state coverage is provided in Appendix D.1. In contrast, SISL demonstrates superior robustness across all evaluated environments, consistently outperforming baselines at varying noise levels. For example, in the Kitchen environment, SISL maintains strong performance under significant noise by effectively refining useful skills, while in Maze2D, higher noise levels lead to the improvement of diverse skills, achieving perfect task completion when $\sigma=1.5$. These results highlight SISL’s ability to discover improved behavior and refine robust skills, significantly enhancing meta-RL performance. Moreover, SISL excels with both noisy and expert data, achieving superior test performance by learning more effective skills.